three times the greater of two consecutive odd integers is five less than four times the smaller. find the two numbers.
Answer by nerdybill(7384)

three times the greater of two consecutive odd integers is five less than four times the smaller. find the two numbers.
.
Let x = smaller consecutive odd integers
then
x+2 = larger consecutive odd integers
.
3(x+2) = 4x-5
3x+6 = 4x-5
6 = x-5
11 = x (smaller integer)
.
larger:
x+2 = 11+2 = 13 (larger integer)